OpenStudy (luigi0210):

Can someone explain this to me? Derivative of an inverse function:

OpenStudy (luigi0210):

Find \(f^{-1}'\) (d) for the given \(f\) and \(d\) \(f(x)=\sqrt{3x+1}\), \(d=1\)

OpenStudy (amistre64):

let y = f(x), then f(y)=x is its inverse dy/dx = f' and dx/dy = (f^-1)' which is just 1/(dy/dx)

OpenStudy (amistre64):

i dont know what (d) is in this context tho ... maybe the first derivative? or maybe x=d=1??

OpenStudy (luigi0210):

I'm given this too: If \(c\) is a value of \(x\), in the interval \([a, b]\) and \(f(c)=d\), then \(f^{-1}'(d)\)=\(\frac{1}{f'(c)}\) I'm assuming you solve for d.. our teacher threw this at us just last night.

OpenStudy (amistre64):

then d is a value for x ... and yeah, that states the same thing invertible functions have reciprocal rates of change.

OpenStudy (amistre64):

of f and f^-1 are inverses of each other such that: f(f^-1(x)) = f^-1(f(x)) then their derivatives have the relationship that:\[(f^{-1})'=\frac 1{f'}\]

OpenStudy (amistre64):

forgot the =x in that lol if ..f(f^-1(x)) = f^-1(f(x)) = x

OpenStudy (amistre64):

in this case: what is the derivative of: sqrt(3x+1) ?

sammixboo (sammixboo):

Good job, Amistre